OpenStudy (anonymous):

A soccer ball is kicked in space, where we assume there is no friction or other forces acting on the ball, at 13 m/s. How much force is needed to keep the ball moving?

OpenStudy (anonymous):

\[S=\frac{1}{2}at^2+vt+s_0\] \[S=\frac{1}{2}at^2+vt\] \[V=at+V_0\] assuming the ball was sitting in space at rest \[V=at=13\] \[\frac{13}{a}=t\]

OpenStudy (anonymous):

None. since there is no other forces impeding its movement, and since it is moving with uniform velocity, then it will, by Newton's first law, continue to move.

OpenStudy (anonymous):

I thought the question asked what was the force upon the ball -.-

OpenStudy (anonymous):

yes. the question asks what is the force on the ball that keeps it moving.

OpenStudy (anonymous):

In the case of the force required to get the ball to 13 m/s, it also depends on how long you exert the force and the mass of the ball. If you apply 10000N but only for a microsecond, you won't gain much velocity. The relation between force and acceleration is \[F = m*a = m * \frac{ \Delta v }{ \Delta t }\] Then, to gain a speed of 13 m/s, the relation is: \[13 = \frac{ F * \Delta t }{ m }\]